OpenStudy (anonymous):

What is the 19th term of the arithmetic sequence 14, 22, 30, … ?

OpenStudy (anonymous):

Here first term is 14 And common difference is : 22 - 14 = 8 So use: \[\large t_n = a + (n-1)d\] Put n = 19 here and a = 14 and d = 8.. and find the 19th term..